OpenStudy (anonymous):

what is the index of a square root mean in lay peoples terms. I know it is the # on the outside of the radical symbol, but what does that translate into words?

OpenStudy (anonymous):

the index of a square root is 2

OpenStudy (anonymous):

If you have a radical symbol with a kth index, how do you explain what the kth index means. Like with an exponent 5^4=5x5x5x5. can you explain it in similar terms?

OpenStudy (anonymous):

\[^k\sqrt{n}=n^{\frac{1}{k}}\]